# Calibrated Physics-Informed Uncertainty Quantification

### v1

@@ 5.5 Magnetic Equilibrium in a Tokamak
Tokamaks confine plasma within a toroidal vessel using magnetic fields to achieve nuclear fusion. The plasma, at high temperatures, is contained by magnetic fields that counterbalance its kinetic pressure. This equilibrium state, a function of magnetic coil configurations and plasma parameters, is governed by the Grad-Shafranov (GS) equation (Somov, 2012):
@@
where $\psi$ represents the poloidal magnetic flux, $p$ the kinetic pressure, $F=rB$ the toroidal magnetic field, and $\mu_{0}$ the magnetic permeability. While traditional numerical solvers like EFIT++ and FreeGSNKE (Lao et al., 1985; Amorisco et al., 2024) are used for equilibrium reconstruction, their computational cost has motivated neural network alternatives (Joung et al., 2023; Jang et al., 2024). However, these surrogate models lack uncertainty quantification capabilities.
Figure 8: Grad-Shafranov: The PRE for a specific poloidal field coil configuration is indicated on the left, and the lower and upper bars for $50\%$ are displayed adjacent to it. Aside from guaranteeing coverage, the PRE-CP framework allows us to discard physically inconsistent equilibria predicted by the surrogate model.
We implement an auto-encoder that maps poloidal magnetic flux across the poloidal cross-section for given tokamak architectures, conditioned on poloidal field coil locations under constant plasma current. While this accelerates simulation by 10000x, it lacks physical guarantees. By incorporating Equation 18 within the PRE-CP framework, we identify physically stable equilibria and obtain statistically valid error bounds. Figure 8 shows the PRE over a surrogate model prediction with lower and upper error bars for $50\%$ coverage. Further details about the problem setting and the model can be found in Appendix N.
@@ 6 Discussion
If “All models are wrong, but some are useful” (Box, 1976), through this work, we explore a novel framework for providing data-free, model and domain agnostic measure of usefulness of neural PDEs. We deploy a principled method of evaluating the accuracy of the solution, i.e. its (calibrated) obedience to the known physics of the system under study. As opposed to other methods of UQ for neural PDEs, our method is physics-informed, allowing us to study the physical inconsistencies of the model predictions with coverage guarantees provided by conformal prediction. This calibration procedure is not limited to PDE modelling but may apply to ODEs or any other scenario where the model outputs can be framed as a residual. We conclude with a discussion of the strengths, limitations and potential improvements.
@@
PRE estimates the violation of conservation laws in neural PDE predictions, guaranteed error bounds over the physics deviation. This post-hoc uncertainty quantification is model- and physics-agnostic, scaling linearly with model complexity and quasi-linearly with PDE complexity due to the additive nature of differential operators. Our framework reformulates CP to be data-free, expressing model inaccuracy solely through PRE, not requiring a labelled dataset. This approach reduces calibration costs and loosens exchangeability restrictions as we can modify the calibration and, hence, the prediction domain by simply reformulating the PRE accordingly. The PRE formulation (Section 4, Appendix B) yields input-independent prediction sets, allowing for the identification of weak predictions within single simulations (marginal-CP) and across multiple predictions (joint-CP). The latter enables a rejection criterion for a set of predictions potentially serving as an active-learning pipeline for neural PDE solvers (Musekamp et al., 2024). PRE-CP provides guaranteed coverage irrespective of the model, chosen discretisation, or the PDE of interest; however, the width of the error bar indicates quantitative features in the model quality. A well-trained model will exhibit tighter error bars as opposed to a poorer fit model, as is demonstrated in Appendix I.
@@
Our method’s coverage bounds exist in the PDE residual space rather than the Euclidean space of physical variables. Transforming to physical space involves challenging set propagation through integral operations, which may require approximations (Teng et al., 2023) or expensive Monte Carlo sampling (Andrieu et al., 2003). The data-free approach lacks a grounding target for calibration, though we argue that a large sample of model outputs provides a statistically significant overview of uncertainty. The sampling cost from the neural-PDE solver for calibration involves intensive gradient evaluations. PRE estimation using finite-difference stencils also introduces the errors associated with Taylor expansion. The current formulation is limited to regular grids with fixed spacing, though extensions to unstructured grids via graph convolutions are possible (Eliasof & Treister, 2020).
@@ 7 Conclusion
We address the problem of reliability of neural-PDE solvers by proposing PRE-CP, a novel conformal prediction framework. Our method provides guaranteed and physics-informed uncertainty estimates for each cell within a prediction, identifying erroneous regions while discerning physically inconsistent predictions across the entire spatio-temporal domain. Our work enhances the reliability of neural PDE solvers, potentially broadening their applicability in science and engineering domains where robust uncertainty quantification is crucial.
@@ Appendix B PRE: Score Function and Prediction Sets
$\mathbb{C}^{\alpha}_{AER}(X_{n+1})=[\hat{f}(X_{n+1})-\hat{q}^{\alpha},\hat{f}(X_{n+1})+\hat{q}^{\alpha}],$
$\mathbb{C}^{\alpha}_{STD}(X_{n+1})=[\hat{f}_{\mu}(X_{n+1})-\hat{q}^{\alpha}\hat{f}_{\sigma}(X_{n+1}),\hat{f}_{\mu}(X_{n+1})+\hat{q}^{\alpha}\hat{f}_{\sigma}(X_{n+1})].$
@@
Appendix C Comparison to Other UQ Methods
Within this section, we compare our method (PRE-CP) with other methods of providing uncertainty estimation neural-PDEs. We compare various Bayesian methods (MC Dropout, Deep Ensembles, Bayesian Neural Networks, Stochastic Weighted Averaging) along with standard inductive conformal prediction to our method and demonstrate that our method is capable of providing valid coverage guarantees for both in and out-of-distribution testing without a significant increase in inference times. Table 2 provides a qualitative comparison of our method against the other benchmarks and highlights that our method is data-free, requires no modification or sampling and provides guaranteed coverage in a physics-informed manner. For the sake of simplicity, we confine our comparison studies in Table 3, 4, and 5 to the Wave, Navier-Stokes and Magnetohydrodynamic equations. The index for the following tables is given below.
MC Dropout
Deep Ensemble
BNN
SWA-G
CP-AER
CP-PRE (Ours)
[…]
in-distribution out-distribution Time
UQ L2 Coverage L2 Coverage Train (hr) Eval (s)
Deterministic 1.77e-05 $\pm$ 3.69e-07 - 2.46e-03 $\pm$ 2.00e-05 - 0:38 22
CP-AER 1.76e-05 $\pm$ 4.40e-07 95.70 $\pm$ 0.21 2.46e-03 $\pm$ 1.41e-05 95.59 $\pm$ 0.14 0:38 23
CP-PRE (Ours) 1.78e-05 $\pm$ 4.61e-07 95.52 $\pm$ 0.21 2.46e-03 $\pm$ 1.25e-05 95.39 $\pm$ 0.12 0:38 23
[…]
in-distribution out-distribution Time
UQ L2 Coverage L2 Coverage Train (hr) Eval (s)
Deterministic 1.05e-04 $\pm$ 6.91e-06 - 3.67e-03 $\pm$ 5.30e-05 - 3:22 25
CP-AER 1.05e-04 $\pm$ 6.58e-06 95.56 $\pm$ 0.40 3.66e-03 $\pm$ 2.81e-05 95.54 $\pm$ 0.15 3:22 26
CP-PRE (Ours) 1.07e-04 $\pm$ 5.18e-06 95.44 $\pm$ 0.22 3.70e-03 $\pm$ 4.23e-05 95.57 $\pm$ 0.14 3:22 34
[…]
in-distribution out-distribution Time
UQ L2 Coverage L2 Coverage Train (hr) Eval (s)
Deterministic 2.20e-03 $\pm$ 5.20e-03 - 4.71e-02 $\pm$ 1.06e-03 - 5:00 40
CP-AER 2.20e-03 $\pm$ 4.38e-05 95.61 $\pm$ 0.26 4.69e-02 $\pm$ 8.18e-04 95.60 $\pm$ 0.27 5:00 42
CP-PRE (Ours) 2.20e-03 $\pm$ 4.96e-03 95.54 $\pm$ 0.18 4.71e-02 $\pm$ 1.06e-03 95.67 $\pm$ 0.22 5:00 82
Table Index
Deterministic: Vanilla FNO (Li et al., 2021)
MC Dropout: FNO with Dropout (Gal & Ghahramani, 2016)
Deep Ensemble: Ensemble of FNOs (Lakshminarayanan et al., 2017)
SWA-G: Stochastic Weighted Averaging - Gaussian (Maddox et al., 2019)
in-distribution: Model evaluated on initial states sampled from the same parameter range (as given in the appendix) of the initial condition as used in the training data.
out-distribution: Model evaluated on initial states sampled from a different parameter range of the initial conditions as used in the training data.
L2: L2 norm of the model output with the ground truth in the normalised domain.
Coverage: Percentage coverage of the model outputs within the estimated error bounds
Train Time: Training time on a single A100 GPU.
Eval. Time: Evaluation time on a single A100 GPU.
@@ Appendix D ConvOperator: Convolutional Kernels for Gradient Estimation
Within the code base for this paper, we release a utility function that constructs convolutional layers for gradient estimation based on your choice of order of differentiation and Taylor approximation. This allows for the PRE score function to be easily expressed in a single line of code  [^2] The code and associated utility functions can be found in this github repository.
@@
First, we use the ConvOperator class from Utils/ConvOps_2d.py to set up the convolutional layer with kernels taken from the appropriate finite difference stencils:
@@ D.1 Impact of Discretisation
As demonstrated in (Bartolucci et al., 2023), the discretisation of the inputs and hence model outputs plays an important role in the accuracy of the neural-PDE solvers. Though the neural operators are constructed for discretisation-invariant behaviour due to the band-limited nature of the functions, they often exhibit discretisation-convergent behaviour rather than be fully discretisation-invariant. This is of particular importance in the temporal dimensions as these neural-PDE models utilise a discrete, autoregressive based time-stepping and is baked into the model within its training regime (McCabe et al., 2023). Due to lack of control in the discretisation within the temporal domain $(dt)$, the PRE estimates tend to have higher numerical errors as well. In fig. 9, we visualise the evaluation of finite difference in 2D+time as a 3D convolution. The finite difference stencil i.e. the convolutional kernel has a unit discretisation of $dx$, $dy$ and $dt$ associated with the problem and is applied over the signal i.e. the output from the neural-PDE $u$ spanning the domain $x,y,t$, where $x\in[0,X],\;y\in[0,Y],\;t\in[0,T]$.
@@
(a) PRE-CP over coarser discretisation
(b) PRE-CP over finer discretisation
@@
Figure 10: PRE-CP provides guaranteed coverage irrespective of the discretisation associated with the model outputs., however, the width of the obtained coverage bounds indicates the discretisation error associated with the gradient estimation. Coverage taken for $\alpha=0.1\sim 90\%$ coverage.
@@ Appendix E Initial and Boundary Conditions
By deploying the eqn 20 as the PRE across the boundary, we can obtain error bars over the boundary conditions as well. Within fig. 11, we demonstrate the error bars obtained by using the boundary conditions as the PRE nonconformity scores for the Navier-Stokes equations.
@@ F.1 Advection Equation
The state variable of interest $u$ is bounded within the domain $x\in[0,2],\;t\in[0,0.5]$ and moves within the domain at a constant velocity $v$. Data generation is performed by solving Equation 21 using a Crank-Nicolson method (Crank & Nicolson, 1947). Data is sampled using a parameterised initial condition that characterises the amplitude and position of the Gaussian field. Generated data is used to train a 1D FNO that takes in the initial condition and autoregressively with a step size of 1, learns to map the next 10 time frames. A reproducible script is attached to the supplementary material.

### v2

@@ 5.5 Magnetic Equilibrium in a Tokamak
Tokamaks confine plasma within a toroidal vessel using magnetic fields to achieve nuclear fusion. The plasma at high temperatures is contained by magnetic fields that counterbalance its kinetic pressure. This equilibrium state, a function of magnetic coil configurations and plasma parameters, is governed by the Grad-Shafranov (GS) equation (Somov, 2012)
@@
where $\psi$ represents the poloidal magnetic flux, $p$ the kinetic pressure, $F=rB$ the toroidal magnetic field, and $\mu_{0}$ the magnetic permeability given in a 2D toroidal coordinate system characterised by $r$ and $z$. While traditional numerical solvers like EFIT++ and FreeGSNKE (Lao et al., 1985; Amorisco et al., 2024) are used for equilibrium reconstruction, their computational cost has motivated neural network alternatives (Joung et al., 2023; Jang et al., 2024). However, these surrogate models lack UQ capabilities, making their deployment within the control room challenging.
Figure 6: Grad-Shafranov: The PRE for a specific poloidal field coil configuration is indicated on the left, and the lower and upper bars for $50\%$ are displayed adjacent to it. Aside from guaranteeing coverage, the CP-PRE framework allows us to discard physically inconsistent equilibria predicted by the surrogate model.
We implement an auto-encoder that maps poloidal magnetic flux across the poloidal cross-section for given tokamak architectures, conditioned on poloidal field coil locations under constant plasma current. While this accelerates simulation by 4 orders of magnitude, it lacks physical guarantees. By incorporating Equation 19 within the CP-PRE framework, we identify physically stable equilibria and obtain statistically valid error bounds. Figure 6 shows the PRE over a surrogate model prediction with lower and upper error bars for $50\%$ coverage. Further details about the problem setting and the model can be found in Appendix N.
6 Extensions of CP-PRE beyond PDEs
Our framework applies to any prediction case where the forward problem can be formulated as a residual with an equality constraint (Appendices A and B). The framework works in any scenario where the forward model can be expressed in the standard canonical form
$Ax-b=0,$ (20)
where $x$ is the model prediction, $A$ is a differential or algebraic operator governing the dynamics, and $b$ is a non-homogeneous term such as a function or a constant.
This extends our applications beyond PDEs to ODEs and algebraic equations found in control problems (Jiang & Jiang, 2014), chemical reactions (Thöni et al., 2025), biological systems (Wang et al., 2019), and financial scenarios (Liu et al., 2019).
We deliberately focus on PDEs as they represent the most comprehensive and challenging case—multi-dimensional domains with complex spatio-temporal dependencies and unique computational challenges. Success with PDEs implicitly validates applicability to simpler systems.
@@ 7 Discussion
If “All models are wrong, but some are useful” (Box, 1976), through this work, we explore a novel framework for providing a measure of usefulness of neural PDEs. We deploy a principled method of evaluating the accuracy of the solution, i.e. its (calibrated) obedience to the known physics of the system under study. As opposed to other methods of UQ for neural PDEs, our method has the unique advantage of being physics-informed. This allows us to study the physical inconsistencies of the model predictions with coverage guarantees provided by conformal prediction. We conclude with a discussion of the strengths, limitations and potential improvements.
@@
PRE estimates the violation of conservation laws in neural PDE predictions, guaranteed error bounds over the physics deviation. This post-hoc uncertainty quantification is model- and physics-agnostic, scaling linearly with model complexity and quasi-linearly with PDE complexity due to the additive nature of differential operators. Our framework reformulates CP to be data-free, expressing model inaccuracy solely through PRE, not requiring a labelled dataset. This approach reduces calibration costs and loosens exchangeability restrictions as we can modify the calibration and, hence, the prediction domain by simply reformulating the PRE accordingly. The PRE formulation (Section 4, Appendix B) yields input-independent prediction sets, allowing for the identification of weak prediction regions within single simulations (marginal-CP) and across multiple predictions (joint-CP). The latter enables a rejection criterion for a set of predictions, potentially serving as an active-learning pipeline for neural PDE solvers (Musekamp et al., 2025). CP-PRE provides guaranteed coverage irrespective of the model, chosen discretisation, or the PDE of interest; however, the width of the error bar indicates quantitative features in the model quality. A well-trained model will exhibit tighter error bars as opposed to a poorer fit model; see Appendix I.
@@
Our method’s coverage bounds exist in the PDE residual space rather than the Euclidean space of physical variables. Transforming to physical space involves challenging set propagation through integral operations, which may require approximations (Teng et al., 2023) or expensive Monte Carlo sampling (Andrieu et al., 2003). The data-free approach lacks a grounding target for calibration, though we argue that a large sample of model outputs provides a statistically significant overview of uncertainty. The sampling cost from the neural-PDE solver for calibration involves intensive gradient evaluations. PRE estimation using finite-difference stencils also introduces the errors associated with Taylor expansion. The current formulation is limited to regular grids with fixed spacing, though extensions to unstructured grids via graph convolutions are possible (Eliasof & Treister, 2020). CP-PRE does not help differentiate between aleatoric and epistemic uncertainty. It aligns with conformal prediction’s characterisation of predictive uncertainty. From one perspective, this could be viewed as aleatoric uncertainty since we construct confidence intervals relative to a specific probability distribution (the distribution from which calibration data, i.e. initial conditions/PDE coefficients, are sampled). Alternatively, it could be considered epistemic uncertainty since we model the neural network’s error through confidence intervals (typically used for unknown but fixed quantities). While we believe the latter interpretation is more appropriate, we acknowledge that the traditional aleatoric/epistemic dichotomy may not be directly applicable to our framework. This distinction is most valuable when both uncertainty types coexist and require separate treatment (Ferson & Ginzburg, 1996).
@@ 8 Conclusion
We address the problem of the reliability of neural-PDE solvers by proposing CP-PRE, a novel conformal prediction framework. Our method provides guaranteed and physics-informed uncertainty estimates for each cell within a prediction, identifying erroneous regions while discerning physically inconsistent predictions across the entire spatio-temporal domain. Our work enhances the reliability of neural PDE solvers, potentially broadening their applicability in science and engineering domains where robust uncertainty quantification is crucial.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix B PRE: Score Function and Prediction Sets
$\mathbb{C}^{\alpha}_{AER}(X_{n+1})$ $=[\hat{f}(X_{n+1})-\hat{q}^{\alpha},\hat{f}(X_{n+1})+\hat{q}^{\alpha}],$ (23)
$\mathbb{C}^{\alpha}_{STD}(X_{n+1})$ $=[\hat{f}_{\mu}(X_{n+1})-\hat{q}^{\alpha}\hat{f}_{\sigma}(X_{n+1}),$
$\quad\hat{f}_{\mu}(X_{n+1})+\hat{q}^{\alpha}\hat{f}_{\sigma}(X_{n+1})].$ (24)
@@
Appendix C Algorithmic Procedure
  1. 1.
    Set up the Neural PDE Solver
    1. (a)
      Define the PDE system of interest with its governing equations in a numerical solver
    2. (b)
      Train a neural network (e.g., Fourier Neural Operator) to approximate solutions to the PDE
    3. (c)
      Ensure the model can make predictions on new initial conditions / PDE coefficients
  2. 2.
    Define the Physics Residual Error (PRE)
    1. (a)
      For a given PDE case with operator D, the physics residual is defined as $|D(\hat{f}(X))|$, where $\hat{f}$ is the neural PDE prediction
    2. (b)
      Create a differential operator using finite difference stencils implemented as convolutional kernels. This operator evaluates how well the predicted solution satisfies the underlying physics
  3. 3.
    Generate Calibration Set
    1. (a)
      Sample a set of initial conditions ${X_{1},X_{2},...,X_{n}}$ from the domain of interest
    2. (b)
      Run the neural PDE solver on these initial conditions to get predictions ${\hat{f}(X_{1}),\hat{f}(X_{2}),...,\hat{f}(X_{n})}$
    3. (c)
      Compute the PRE nonconformity scores ${|D(\hat{f}(X_{1}))|,|D(\hat{f}(X_{2}))|,...,|D(\hat{f}(X_{n}))|}$ for each prediction
  4. 4.
    Calibration Process
    1. (a)
      For a desired confidence level 1-$\alpha$ (e.g., $90\%$ for $\alpha=0.1$), compute the empirical quantile:
    2. (b)
      $\hat{q}^{\alpha}=\lceil(n+1)(1-\alpha)\rceil/n$ quantile of ${|D(\hat{f}(X_{1}))|,|D(\hat{f}(X_{2}))|,....,|D(\hat{f}(X_{n}))|}$
    3. (c)
      This quantile represents the threshold for the physics residual that will provide the desired coverage
  5. 5.
    Apply to New Predictions
    1. (a)
      For a new initial condition $X_{n+1}$, obtain the neural PDE prediction $\hat{f}(X_{n+1})$
    2. (b)
      Compute the physics residual $|D(\hat{f}(X_{n+1}))|$
    3. (c)
      The prediction set is defined as $\mathbb{C}^{\alpha}=[-\hat{q}^{\alpha},\hat{q}^{\alpha}]$
    4. (d)
      If $|D(\hat{f}(X_{n+1}))|$ falls within $\mathbb{C}^{\alpha}$, the prediction is considered valid at the $1-\alpha$ confidence level
  6. 6.
    Interpretation of Results
    1. (a)
      For marginal conformal prediction: Apply steps 1-5 cell-wise across the spatial-temporal domain to get localized error bounds
    2. (b)
      For joint conformal prediction: Calculate the supremum of the normalized residuals across the entire domain to obtain global error bounds
    3. (c)
      The width of the error bounds indicates the model’s physical consistency - tighter bounds suggest better alignment with the physics
  7. 7.
    Decision Framework (Optional)
    1. (a)
      Accept predictions where the physics residual falls within the calibrated bounds
    2. (b)
      Reject and flag predictions where the residual exceeds the bounds, potentially routing these cases to traditional numerical solvers
    3. (c)
      This approach guarantees that, with probability at least $1-\alpha$, the physics residual of new predictions will fall within the computed bounds, ensuring that the model maintains physical consistency while providing faster solutions than traditional numerical methods.
@@ Appendix D ConvOperator: Convolutional Kernels for Gradient Estimation
Within the code base for this paper, we release a utility function that constructs convolutional layers for gradient estimation based on your choice of order of differentiation and Taylor approximation. This allows for the PRE score function to be easily expressed in a single line of code  [^2] The code and associated utility functions can be found in: https://github.com/gitvicky/CP-PRE
@@
First, we use the ConvOperator class from Utils/ConvOps_2d.py to set up the convolutional layer with kernels  [^3] Convolution functions can be set as cross-correlations as it is default in the PyTorch framework, or they could be set up as convolutions by obtaining the complex conjugate in the frequency space.(Jordà et al., 2019) taken from the appropriate finite difference stencils:
@@ D.1 Impact of Discretisation
As demonstrated in (Bartolucci et al., 2023), the discretisation of the inputs and hence model outputs plays an important role in the accuracy of the neural-PDE solvers. Though the neural operators are constructed for discretisation-invariant behaviour due to the band-limited nature of the functions, they often exhibit discretisation-convergent behaviour as opposed to discretisation-invariance. This is of particular importance in the temporal dimensions as these neural-PDE models utilise a discrete, autoregressive based time-stepping, baked into the model within its training regime (McCabe et al., 2023). The lack of control in teh temporal discretisation $(dt)$, leads to higher numerical errors within the the PRE estimates. In fig. 7, we visualise the evaluation of finite difference in 2D+time as a 3D convolution. The finite difference stencil i.e. the convolutional kernel has a unit discretisation of $dx$, $dy$ and $dt$ associated with the problem and is applied over the signal i.e. the output from the neural-PDE $u$ spanning the domain $x,y,t$, where $x\in[0,X],\;y\in[0,Y],\;t\in[0,T]$. Though the discretisation deployed within CP-PRE stems from the neural PDE solver, it consistently delivers guaranteed coverage regardless of resolution. Even with coarser discretisation, CP-PRE remains valuable as it allows for: statistical identification of poorer fit regions (marginal formulation), highlights physically inconsistent predictions (joint formulation) and enables relative assessment of physical inconsistencies across predictions. While residuals may be inflated with coarser discretisation, the corresponding bounds reflect this inflation, preserving the relative information about physical inconsistency across a series of predictions.
@@
(a) CP-PRE over coarser discretisation
(b) CP-PRE over finer discretisation
@@
Figure 8: CP-PRE provides guaranteed coverage irrespective of the discretisation associated with the model outputs., however, the width of the obtained coverage bounds indicates the discretisation error associated with the gradient estimation. Coverage taken for $\alpha=0.1\sim 90\%$ coverage.
@@ Appendix E Initial and Boundary Conditions
By deploying the eqn 26 as the PRE across the boundary, we can obtain error bars over the boundary conditions as well. Within fig. 9, we demonstrate the error bars obtained by using the boundary conditions as the PRE nonconformity scores for the Navier-Stokes equations.
@@ F.1 Advection Equation
The state variable of interest $u$ is bounded within the domain $x\in[0,2],\;t\in[0,0.5]$ and moves within the domain at a constant velocity $v$. Each trajectory is generated by solving Equation 27 using a Crank-Nicolson method (Crank & Nicolson, 1947). Data is sampled using a parameterised initial condition that characterises the amplitude and position of the Gaussian field. Generated data is used to train a 1D FNO that takes in the initial condition and autoregressively with a step size of 1, learns to map the next 10 time frames as outlined in Equation 28. A reproducible script is attached to the supplementary material.
$u^{n+1}=f(u^{n}),$ (28)
$f$ represents the FNO and $u^{n}$ and $u^{n+1}$ represents the current (input) and the predicted future state of the system (output).